This page gives a list of the more common errors experienced with IRCAM. A Diagnostics Guide is also available, for internal use only, from the UKIRT reference page . If all else fails, the observers should exit ORACOM, the TSS should run down the camera, log off IRTCON, and then log on and bring up the instrument again. This often clears up and restarts troubled processes. The errors are listed in approximate order of how commonly they are seen at the telescope.
Under the ORAC system the first indication of a BDS error will be the following error message on the observer’s OM screen:
The will also be error messages in the xterm from which oracom was launched on the observers side:
To recover, the TSS goes up the menus to the “Engineering” menu, go into “ALICE Engineering”, and “Kill and Reload OCCAM”. The pinkish ALICE screen will disappear and return. You can then hit all the OK buttons on any OM error messages on both the observers and the TSS side. The observer should then hit stopASAP, move the highlight back up a line to the “set object” or “set dark” preceding the failed observation, then hit run-from-highlight, and the missing observation will be repeated. (If you simply hit continue you will lose the last data frame and frame number. You can recover in the DR by using the “- skip” option. ) BE CAREFUL: about half the time the OM stops saving data after a BDS error so watch for this and if there’s a problem run the OM down and up.
BDS errors tend to occur when you are taking data at fast rates. If your exposures are 1 second or less, try using STARE instead of ND_STARE. Or cut down on the number of coadds.
Filter Detent Errors
It is common for the camera to have to hunt for the filter position, and you will see a message something like:
Filter found after 20 steps
Ignore searches at the level of about 20 steps and continue taking data. However if the filter is still not found after a couple of searches, or if the number of steps is large, then go to the “Engineering” menu and datum the filter wheels.
Have patience if you see the exposure countdown go to negative counts of a few seconds, as there are overheads associated with every read, but if it gets to say -50 then you’re probably in trouble. You can try the pf1 or pf4 keys to get out, but usually the only way out is a control-c. To ensure all processes are properly killed and restarted, log completely off of IRTCON, log on, and bring back up IRCAM3_DEV.
On occasion the IEEE Bus process that controls the camera motors has to be restarted. You will see an error something like:
High Dark Current
On occasion the first measurement of dark current gives a slightly high value. Run array tests again. If the value is still high check the instrument temperatures and if they are high call Mike Wagner or Sandy Leggett.
Frames Not Offsetting
On occasion the instrument and telescope lose communication and you will see that your jittered frames are not in fact offset. Check the OM screen says “Telescope: connected”, if it does check the instrument engineering screens for communication flags. If the OM side says disconnected run the OM down and up.